If the product of a number and -5 is reduced by 7, the result is 15 less than three times the opposite of the number. Find the number.
Answer by ankor@dixie-net.com(22740)

If the product of a number and -5 is reduced by 7, the result is 15 less than three times the opposite of the number. Find the number.
:
let n = the number
then
-n = the opposite of the number
:
-5n - 7 = 3(-n) - 15
-5n + 3n = -15 + 7
-2n = -8
n = -8/-2
n = +4
:
:
See if that works
-5(4) - 7 = 3(-4) - 15
-20 - 7 = -12 - 15